# Uncovering the role of semantic and acoustic cues in normal and dichotic listening

### v1

@@ IV-C Dichotic Listening: Speech Envelope vs Semantics
The results, presented in Table II, demonstrate that the stimulus-response model using semantic features as input outperformed the model utilizing speech envelope features (Wilcoxon signed rank test, ). This indicates that semantic features are better entrained in the EEG response recorded during the dichotic listening experiment. The multi-modal network achieves an average accuracy of %, compared to % with the speech feature-based network. This significant performance improvement (Wilcoxon signed rank test, ) indicates that the EEG signal jointly encodes the semantic and acoustic content of the stimulus. Furthermore, this MM performance signifies that the proposed multi-modal model is able to correctly identify the intended channel of attention in more than % of the segments, a capability that can potentially influence BCI application design. The findings also suggest that semantics may hold higher significance than acoustic traits for speech perception under complex listening conditions.
@@ IV-D Region-wise analysis
The region-wise analysis findings shown in Table VI, revealed that the temporal region exhibited the highest performance in the mismatch (MM) classification task, followed by the occipital, parietal, frontal, and central regions. The performance improvement is also observed to be consistent for semantic and acoustic features. This observed pattern suggests that certain brain regions may play more critical roles in processing dichotic listening stimuli, particularly in tasks involving attention and speech processing.
IV-E Attention to right ear versus left ear in dichotic listening
Additionally, the results presented in Table VII show that subjects exhibited significantly greater attention to the stimulus presented in the right ear (S2) compared to the left ear (S1). This supports the phenomenon of right ear dominance[43] in dichotic listening tasks, where individuals typically pay greater attention to stimuli delivered to the right ear.
@@ V Conclusion
In this study, we compared EEG-stimulus models under both natural speech and dichotic listening conditions, aiming to assess the relative contribution of acoustic and semantic features in different listening scenarios. Our findings reveal that the significance of textual and acoustic content is similar during natural speech listening. However, in the dichotic speech listening task, EEG signals demonstrated a higher match-mismatch performance for textual data compared to the audio signal. This suggests a more pronounced encoding of higher-level semantic information over acoustic envelope information in dichotic listening condition.
The study also investigated the role of word boundary information in the MM task performance. It is seen that the word boundary information plays a pivotal role in speech perception during the dichotic listening task. Notably, not all boundaries seem to be critical as in the case of natural speech. Further, the study added further evidence to right ear advantage observed in dichotic listening through the match-mismatch classification experiments.

### v2

@@ IV-C Dichotic Listening: Speech Envelope vs Semantics
During dichotic listening, the human brain exhibits a prioritization of perceiving the semantics over the acoustic characteristics of the attended speech (Wilcoxon signed-rank test, ). This indicates that semantic features are better entrained in the EEG response recorded during the dichotic listening experiment. The multimodal network achieves an average accuracy of %, compared to % with the speech feature-based network. This significant performance improvement (Wilcoxon signed-rank test, ) indicates that the EEG signal jointly encodes the semantic and acoustic content of the stimulus. The findings also suggest that semantics may hold higher significance than acoustic traits for speech perception under complex listening conditions. Furthermore, this MM performance signifies that the proposed multimodal model is able to identify the intended channel of attention in % of the segments, a capability that can potentially influence in cognitive design of hearing aids.
@@ IV-D Region-wise analysis
The region-wise analysis, shown in Table IV, revealed that the temporal region exhibited the highest performance in the mismatch (MM) classification task, followed by the parietal, occipital, frontal, and central regions. The performance trend is also observed to be consistent for semantic and acoustic features. The observed pattern (Figure 6) suggests that temporal brain regions may play more critical roles in processing dichotic stimuli, particularly in tasks involving attention and speech processing. These findings are also consistent with those reported by Hausfeld et al. [52] and Ross et al. [53].
IV-E Right ear advantage in dichotic listening
Additionally, the results presented in Table V show that subjects exhibited significantly higher attention to the stimulus presented in the right ear (S2) compared to the left ear (S1) (Wilcoxon rank-sum test ). This supports the phenomenon of right ear dominance [26], which is often considered as neurophysiological bias especially for speech stimulus in dichotic listening tasks, where individuals typically pay greater attention to stimuli delivered to the right ear. We find these results to align with observations of Tanaka et al. [27] on the asymmetry of hemispheric organization and with Kazimierczak et al. [54] on auditory laterality in brain.
IV-F Limitations and Future Directions
The lack of sharply defined functional boundaries in the brain and the inherent complexity of brain activity limit the precision of analyses based on non-invasive EEG signals. While the proposed architectures for normal and dichotic listening offer valuable insights into the semantic and acoustic components of attention, the current experimental setup does not fully reflect real-world conditions. Factors such as multiple spatially distributed sound sources, listener orientation, and the influence of visual and other sensory cues—each of which might affect auditory attention—are not accounted for. Future work will focus on extending the paradigm to incorporate more ecologically valid scenarios by integrating these real-world variables into stimulus design.
@@ V Conclusion
In this study, we compared Stimulus-EEG match-mismatch models under both natural speech and dichotic listening conditions, aiming to assess the relative contribution of acoustic and semantic features in different listening scenarios. Our findings reveal that the significance of textual and acoustic content is similar during natural speech listening. However, in the dichotic speech listening task, EEG signals demonstrated a higher MM performance for textual data compared to the audio signal. This suggests a more pronounced encoding of higher-level semantic information over acoustic envelope information in dichotic listening condition. The study also investigated the role of word boundary information in the MM task performance. It is seen that the word boundary information plays a pivotal role in speech perception during the dichotic listening task. Further, the study added further evidence to right ear advantage observed in dichotic listening through the MM classification experiments.